The number of typos in a book has a poisson distribution with an average of 5 typos
per 100 pages. What is the probability that you cannot find a typo in 50 pages.
1
Expert's answer
2020-04-28T18:41:39-0400

P(X=x)=(λxe-λ )/x!

λ = 5 typos per 100 pages = 2.5 typos per 50 pages

P=P(X=0)=(2.50*e-2.5)/0!=0.0821